OpenStudy (anonymous):

Challenge: Find the ratio of area of triangle and area of triangle formed using the lengths of the medians.

OpenStudy (anonymous):

Sorry it is not a challenge it took me 2mins to realize the solution

Directrix (directrix):

If we're talking a single triangle, the area would be constant and the ratio 1:1.

OpenStudy (anonymous):

No

OpenStudy (anonymous):

area of the triangle of medians is three-fourths of the area of the given triangle.

OpenStudy (anonymous):

three medians of a triangle divide the triangle into six equal areas.

OpenStudy (anonymous):

That is right aron

OpenStudy (anonymous):

However tell me how did you do it just want to check if we both used the same method

OpenStudy (anonymous):

Yep it's 3/4.

OpenStudy (anonymous):

I used this: http://mathworld.wolfram.com/TriangleMedian.html

OpenStudy (anonymous):

I know a proof using vectors

OpenStudy (anonymous):

Aha, vector would certainly make things much easy.

OpenStudy (anonymous):

It is a 3 line proof And i discovered it myself

OpenStudy (anonymous):

Furthermore, the area of the triangle is : \[1/3\sqrt{(2\sum u ^{2}v ^{2} - \sum u^{4})}\] where u,v and w are the medians .

OpenStudy (anonymous):

@Aron that's a fairly complicated form.